OpenStudy (anonymous):

10x-5y=-20 7x-5y=-26 find the solutions

OpenStudy (anonymous):

use substitution method :)

OpenStudy (anonymous):

since it only 2 equations and 2 variables

OpenStudy (anonymous):

okaii subsitiuion method???

OpenStudy (anonymous):

10x-5y=-20 7x-5y=-26 1st solve for y in eq 1 -5y=-20 -10x y=4+2x now you plug in 4+2x into eq 2 7x-5(4+2x )=-26 solve for x. once you solve for x. x will be x=n and you use that number to plug back into x and solve for y

OpenStudy (anonymous):

ummmmmmmmm examlpe???

OpenStudy (anonymous):

just stack the two on top of each other and you want to cancel either an x or y and then go from there

OpenStudy (anonymous):

o okaii

OpenStudy (anonymous):

mhm its pretty simple youll see i love these types of problems :) its called elimination

OpenStudy (anonymous):

ughh really??? u too ??

OpenStudy (anonymous):

wait wut

OpenStudy (anonymous):

substitution takes longer i like elimantion you cross out and plug back in super simple

OpenStudy (anonymous):

substitution is good for these problems

OpenStudy (anonymous):

eliminatiion is betteerrrr

OpenStudy (anonymous):

not wat i meant but thanks

OpenStudy (anonymous):

okie dokie